A bag contains 3 blue stones, 5 red stones, and 2 white stones. What is the probability of picking a blue stone, keeping it and then picking a white stone?
Answer by stanbon(57203)

A bag contains 3 blue stones, 5 red stones, and 2 white stones.
What is the probability of picking a blue stone, keeping it and then picking a white stone?
------------------
(3/10)(2/9) = 2/30 = 1/15
===============================
Cheers,
Stan H.